# TomoGraphView: 3D Medical Image Classification with Omnidirectional Slice Representations and Graph Neural Networks

### v1

@@ Abstract
The growing number of medical tomography examinations has necessitated the development of automated methods capable of extracting comprehensive imaging features to facilitate downstream tasks such as tumor characterization, while assisting physicians in managing their growing workload. However, 33D medical image classification remains a challenging task due to the complex spatial relationships and long-range dependencies inherent in volumetric data. Training models from scratch suffers from low data regimes, and the absence of 33D large-scale multimodal datasets has limited the development of 33D medical imaging foundation models. Recent studies, however, have highlighted the potential of 22D vision foundation models, originally trained on natural images, as powerful feature extractors for medical image analysis. Despite these advances, existing approaches that apply 22D models to 33D volumes via slice-based decomposition remain suboptimal. Conventional volume slicing strategies, which rely on canonical planes such as axial, sagittal, or coronal, may inadequately capture the spatial extent of target structures when these are misaligned with standardized viewing planes. Furthermore, existing slice-wise aggregation strategies rarely account for preserving the volumetric structure, resulting in a loss of spatial coherence across slices. To overcome these limitations, we propose TomoGraphView, a novel framework that integrates omnidirectional volume slicing with spherical graph-based feature aggregation. Unlike traditional methods, which are restricted to canonical views, our approach samples both canonical and non-canonical cross-sections. These non-canonical views are derived from uniformly distributed points on a sphere, which visually encompasses the 3D volume, thereby producing a richer set of cross-sectional representations. As the spherically distributed viewpoints naturally define a spherical graph topology after triangulation, we allow for the explicit encoding of spatial relationships across views as nodes and corresponding edges in the underlying graph topology, and leverage a graph neural network for spatial-aware feature aggregation. Experiments across six oncology 33D medical image classification datasets demonstrate that omnidirectional volume slicing improves the average performance in Area Under the Receiver Operating Characteristic Curve (AUROC) from $0.7701$ to $0.8154$ compared with traditional slicing approaches relying on canonical view planes. Moreover, we can further improve AUROC performance from $0.8198$ to $0.8372$ by leveraging our proposed graph neural network-based feature aggregation. Notably, TomoGraphView surpasses large-scale pretrained 33D medical imaging models across all datasets and tasks, underscoring its effectiveness as a powerful framework for volumetric analysis and therefore represents a key step toward bridging the gap until fully native 33D foundation models become available in medical image analysis.
@@
[label3]organization=Institute of Machine Learning in Biomedical Imaging, Helmholtz Center Munich, city=Neuherberg, country=Germany
@@
[label4]organization=Institute of Radiation Medicine, Helmholtz Center Munich, city=Neuherberg, country=Germany
@@
[label7]organization=Munich Center for Machine Learning, country=Germany
@@ 1 Introduction
Deep learning methods have emerged as promising tools in medical imaging research for their ability to learn complex, hierarchical feature representations from large imaging datasets. Such approaches enable the detection of pathological processes, supporting applications ranging from early disease detection and accurate diagnosis [rana2023machine], malignancy prediction [prakash2020breast, naser2020brain, navarro2021development], and lesion characterization [herent2019detection, yin2023value, dadoun2022deep], to robust image segmentation [ronneberger2015u, isensee2021nnu, cao2022swin, erdur2025deep] and prognostic as well as predictive modeling [zhu2020application, schulz2021multimodal, spohn2023development, kraus2023dosiomics]. These capabilities not only facilitate personalized treatment strategies but also hold the potential to improve therapeutic effectiveness and patient outcomes [rasuli2025deep]. Driven by technological advancements that have led to a wealth of tomography data, image analysis methods have gained increasing relevance in medical research, demanding automated tools to support clinicians [thrall2018artificial]. To this end, the automated and effective extraction of imaging biomarkers from 33D medical imaging data constitutes a core component that substantially alleviates clinical workload, while also enhancing the efficacy of subsequent downstream tasks [rasuli2025deep].
@@
Generally, a promising direction to extract semantically rich imaging features from 22D images is represented by foundation models, which have already attracted significant attention in deep learning based computer vision applications [radford2021learning, kirillov2023segment, zou2023segment, oquab2023dinov2, fang2024eva, tschannen2025siglip, venkataramanan2025franca, simeoni2025dinov3]. They not only show great promise in their ability to extract comprehensive feature representations from imaging data but also in their easy adaptability to a wide range of downstream tasks, such as segmentation [ayzenberg2024dinov2], classification [veasey2024parameter], or registration [song2024dino]. While applying these powerful models has become standard practice in natural image analysis, the progress in adapting them to 33D medical imaging is proceeding more slowly. A key challenge is the lack of large and publicly available multimodal medical imaging datasets required for training at scale [azad2023foundational]. Following paschali2025foundation [paschali2025foundation], a foundation model in radiology should incorporate both large-scale model architectures and multimodal training data, leverage self-supervised training to reduce the need for expert annotations, and crucially, demonstrate emergent capabilities beyond their training objectives. Despite the progress and rapid development of large-scale pretrained 33D models in medical imaging research, performance levels of radiology foundation models still lag behind their 22D counterparts [van2025foundation]. Regardless of challenges such as the limited availability of large-scale 33D datasets and the computational demands of training 33D models, ongoing research in these areas is likely to lead to substantial breakthroughs and enhanced integration of these models into clinical workflows [azad2023foundational, van2025foundation].
@@
Despite 33D large-scale pretrained models lagging behind the performance levels of their 22D counterparts, recent studies have demonstrated that 22D vision foundation models trained on natural images can be effectively transferred to medical imaging tasks [huix2024natural, liu2025doesdinov3setnew]. In particular, the DINO [caron2021emerging] family of models has shown considerable promise in the medical imaging domain [huang2024comparative, cui2024surgical, baharoon2024evaluatinggeneralpurposevision, perez2025exploring]. However, applying 22D models to 33D input data necessitates slice-based decomposition of the input volumes, resulting in the loss of structural information across slices, as slice-wise representations disregard the three-dimensional spatial context inherent in 33D volumes. A common strategy to combine the resulting slice-wise information extracted by a pretrained encoder is to use a dimension-wise average thereof or a shallow multi-layer perceptron (MLP) [kiechle2024graph, muller2025medical, liu2025doesdinov3setnew]. While this enables end-to-end learning for downstream tasks such as classification, there is no inherent mechanism that explicitly accounts for the relative positions among the slices. To address this limitation, more structured approaches have been explored. Long short-term memory (LSTM)–based aggregation methods [donahue2015long, nguyen2020cnn, lilhore2025hybrid] treat consecutive slices as temporal sequences, thereby capturing inter-slice relationships. Alternatively, transformer-based aggregation [muller2025medical, meneghetti2025end] leverages self-attention mechanisms to model spatial relationships across slices more directly.
@@
With the emergence of graph neural networks (GNNs) [scarselli2008graph], effective modeling of naturally graph-structured data has become possible. GNNs are particularly well-suited for problems involving pairwise interactions or modeling complex relational data by encoding relationships in the graph structure [wu2020comprehensive]. In the context of 33D data, GNNs have demonstrated a strong ability to preserve an underlying spatial structure when the graph topology is based on the relative position of 22D representations [Wei_2020_CVPR, Wei_2023_IEEE]. This capability can be extended to 33D medical data, where GNNs can effectively represent spatial relationships between encoded slices within the input volume [kiechle2024graph, di2025structured].
@@
In this work, we present TomoGraphView, a novel framework designed to preserve the spatial structure of volumetric medical imaging data when decomposed into 22D slices compatible with 22D pretrained encoders. The name reflects its core components: Tomo refers to tomography data, Graph denotes the use of GNN-based feature aggregation, and View signifies the integration of multi-view information from omnidirectional perspectives. TomoGraphView consists of two main components: Omnidirectional volume slicing and subsequent spherical graph-based feature aggregation. In our approach, volumetric data, such as CT or MRI scans, are represented as structured graphs, with each graph corresponding to a single tomography volume. The nodes represent features encoded by a 22D vision foundation model given cross-sectional views, covering both canonical and non-canonical planes. The graph topology is defined by triangulating these viewpoints, thereby approximating a 33D spherical structure. This design enables a more faithful representation of the 33D volume by embedding both the 22D slice information and their spatial relationships within the three-dimensional graph. Edges between nodes are weighted according to their relative distance, allowing the model to efficiently integrate local and global context while explicitly preserving spatial structure. By aggregating node features through a GNN, our framework effectively models 3D context from 2D representations, demonstrating the strength of graph-based feature aggregation for volumetric data analysis. The main contributions and findings of this work can be summarized as follows:
@@
(I) We introduce a novel out-of-plane volume slicing strategy, termed omnidirectional volume slicing, which generalizes beyond conventional axial, coronal, and sagittal slicing and demonstrates superior downstream performance when compared to traditional volume slicing approaches solely relying on canonical axes.
@@
(II) We introduce a spherical mesh-based graph structure derived from our omnidirectional slicing strategy, enabling the preservation of spatial relationships between extracted 2D slices from tomographic data through spherical graph-based feature aggregation.
@@
(III) Extensive experiments on six oncology 33D medical imaging datasets demonstrate that the proposed TomoGraphView framework, integrating omnidirectional volume slicing with spherical graph-based feature aggregation, consistently outperforms baseline slice-wise aggregation methods and large-scale pretrained 33D medical models across all datasets and tasks, establishing TomoGraphView as a powerful framework for 33D medical image classification.
@@ 2.1 Slice-based Methods in 33D Medical Imaging
Two-dimensional slice-based models continue to be an attractive alternative to heavy 33D models, as they are computationally lighter, benefit from the decomposition of 33D volumes into numerous 22D slices, and can leverage powerful pretrained backbones from natural image analysis, which have already demonstrated strong performance levels in medical image classification tasks [huix2024natural, liu2025doesdinov3setnew].
Several studies have demonstrated the effectiveness of 22D approaches. navarro2021development [navarro2021development] fine-tuned a DenseNet161 pretrained on ImageNet using axial T1- and T2-weighted MRI slices of soft-tissue sarcomas for tumor grading. Similarly, ahamed2023slice [ahamed2023slice] trained a ResNet18 on axial PET/CT slices of lymphoma to classify whether a slice intersected tumor tissue. wang2024svfr [wang2024svfr] proposed a two-stage slice-to-volume feature representation framework for Alzheimer’s disease, where informative axial MRI slices were selected and aggregated into volume-level features. In parallel, self-supervised foundation models have also been evaluated for slice-level tasks. huang2024comparative [huang2024comparative] and baharoon2024evaluatinggeneralpurposevision [baharoon2024evaluatinggeneralpurposevision] applied DINOv2 to axial brain MRIs and chest CTs for glioma grading and COVID-19 diagnosis, respectively, while liu2025doesdinov3setnew [liu2025doesdinov3setnew] assessed DINOv3 on non-contrast axial CT slices for detecting 18 clinically significant abnormalities.
Beyond single-plane analysis, multi-plane strategies that combine axial, coronal, and sagittal slices, commonly referred to as 2.52.5D, have been proposed. saint20222 [saint20222] employed an Xception model trained on 2.52.5D representations to predict HPV status in oropharyngeal cancer. Similarly, meneghetti2025end [meneghetti2025end] explored foundation model-based multiple instance learning for head and neck cancer outcomes, comparing single-plane, multi-plane, and 33D sub-volume representations. A related hybrid approach by jang2022m3t [jang2022m3t] first extracted volumetric features using a 33D CNN, followed by multi-plane slicing of the feature tensor and 22D CNN and transformer processing for Alzheimer’s disease classification.
However, none of these methods extract slices from non-canonical orientations, i.e., directions that are not aligned with the standard axes. Consequently, valuable information from structures misaligned with these planes may be omitted. Moreover, while choosing a powerful method for decomposing 33D data into 22D slices is essential, the aggregation of slice-wise features remains equally critical for preserving volumetric context and achieving strong downstream performance, as discussed in the following section.
@@ 2.2 Slice-wise Feature Aggregation in 33D Medical Imaging
Slice-wise feature aggregation represents a promising direction for 3D medical image analysis. By leveraging powerful 22D foundation models as feature encoders, it avoids the high computational cost of native 33D training while allowing a flexible combination of slice embeddings. However, decomposing a 33D volume into 22D slices inherently discards volumetric structure, meaning that spatial dependencies between slices must be explicitly modeled to preserve the structural integrity of the original data.
@@
The simplest approach aggregates slice embeddings into a single volume-level representation using non-learnable operations such as mean pooling. muller2025medical [muller2025medical] employed this strategy by encoding axial slices with DINOv2 and averaging embeddings to classify breast MRI, lung CT, and knee MRI scans across multiple tasks. Similarly, liu2025doesdinov3setnew [liu2025doesdinov3setnew] extracted DINOv3-based embeddings from axial chest CT slices to classify anatomical structures into 18 clinical categories. Although efficient, such mean-pooling schemes ignore inter-slice relationships and therefore likely yield suboptimal performance.
To introduce learnability, several studies replaced pooling with multi-layer perceptron (MLP) aggregation. truong2021transferable [truong2021transferable] fine-tuned multiple self-supervised ImageNet models with an MLP head for various medical classification tasks involving lymph node, fundus, and chest X-ray images. baharoon2024evaluatinggeneralpurposevision [baharoon2024evaluatinggeneralpurposevision] extended this idea to 100 radiology experiments across modalities (X-ray, CT, MRI), while kiechle2024graph [kiechle2024graph] compared MLP-based aggregation across anatomical structures in the MedMNIST3D dataset. Although MLPs provide learnable fusion, they typically neglect slice order and spatial relations unless explicitly conditioned on slice indices [kiechle2024graph].
Recurrent architectures such as long short-term memory (LSTM) networks address this limitation by treating slices as ordered sequences, thus maintaining spatial continuity [donahue2015long]. zhang2019rsanet [zhang2019rsanet] introduced a slice-wise attention network that models MRIs as sequential inputs to capture long-range dependencies for multiple sclerosis lesion segmentation. Similarly, nguyen2020cnn [nguyen2020cnn] combined CNNs with LSTMs for intracranial hemorrhage detection in CT scans, while lilhore2025hybrid [lilhore2025hybrid] integrated bidirectional LSTMs with EfficientNet-B0 for breast cancer classification. miron2021evaluating [miron2021evaluating] also processed volumes slice-by-slice through a CNN, fused feature maps at the channel level, and refined them using a lightweight convolutional head into a prediction.
@@
More recently, transformer-based models have gained traction by leveraging attention to learn inter-slice dependencies. jang2022m3t [jang2022m3t] proposed a multi-plane and multi-slice transformer (M3T) for Alzheimer’s diagnosis, fusing axial, coronal, and sagittal features extracted from a 33D CNN. meneghetti2025end [meneghetti2025end] introduced an end-to-end multiple instance learning (MIL) transformer for predicting clinical outcomes in head and neck cancer, encoding 22D CT slices with a foundation model and aggregating them using transformer-based attention. muller2025medical [muller2025medical] later presented the Medical Slice Transformer (MST), which combines transformer aggregation with 22D self-supervised encoders and demonstrates strong results across breast MRI, lung CT, and knee MRI classification.
@@
A complementary direction represents volumetric data as structured graphs. kiechle2024graph [kiechle2024graph] and di2025structured [di2025structured] modeled tomographic volumes where nodes correspond to slice features and edges encode spatial adjacency. kiechle2024graph [kiechle2024graph] showed that graph representations improve accuracy and runtime over MLP-based aggregation but depend heavily on selecting suitable graph convolution operators and topologies, limiting general applicability. di2025structured [di2025structured] constructed graphs from axial slice stacks to capture both local and global inter-slice dependencies along the axial axis. However, this design remains restricted to a single orientation, omitting valuable coronal and sagittal context. Moreover, evaluation limited to chest CTs leaves the generalizability of this approach across modalities and anatomical sites uncertain.

### v2

@@ Abstract
The sharp rise in medical tomography examinations has created a demand for automated systems that can reliably extract informative features for downstream tasks such as tumor characterization. Although 3D volumes contain richer information than individual slices, effective 3D classification remains difficult: volumetric data encode complex spatial dependencies, and the scarcity of large-scale 3D datasets has constrained progress toward 3D foundation models. As a result, many recent approaches rely on 2D vision foundation models trained on natural images, repurposing them as feature extractors for medical scans with surprisingly strong performance. Despite their practical success, current methods that apply 2D foundation models to 3D scans via slice-based decomposition remain fundamentally limited. Standard slicing along axial, sagittal, and coronal planes often fails to capture the true spatial extent of a structure when its orientation does not align with these canonical views. More critically, most approaches aggregate slice features independently, ignoring the underlying 3D geometry and losing spatial coherence across slices.
To overcome these limitations, we propose TomoGraphView, a novel framework that integrates omnidirectional volume slicing with spherical graph-based feature aggregation. Instead of restricting the model to axial, sagittal, or coronal planes, our method samples both canonical and non-canonical cross-sections generated from uniformly distributed points on a sphere enclosing the volume. Triangulating these viewpoints yields a spherical graph that captures spatial relationships among views, and we use a graph neural network to aggregate their features accordingly. Experiments across six oncology 33D medical image classification datasets demonstrate that omnidirectional volume slicing improves the average performance in Area Under the Receiver Operating Characteristic Curve (AUROC) from $0.7701$ to $0.8154$ compared with traditional slicing approaches relying on canonical view planes. Moreover, we can further improve AUROC performance from $0.8198$ to $0.8372$ by leveraging our proposed graph neural network-based feature aggregation. Notably, TomoGraphView surpasses large-scale pretrained 33D medical imaging models across all datasets and tasks, underscoring its effectiveness as a powerful framework for volumetric analysis and therefore represents a key step toward bridging the gap until fully native 33D foundation models become available in medical image analysis. We provide a user-friendly library for omnidirectional volume slicing at https://pypi.org/project/OmniSlicer.
@@
[label3]organization=Institute of Machine Learning in Biomedical Imaging, Helmholtz Munich, city=Neuherberg, country=Germany
@@
[label4]organization=Institute of Radiation Medicine, Helmholtz Munich, city=Neuherberg, country=Germany
@@
[label7]organization=Munich Center for Machine Learning (MCML), city=Munich, country=Germany
@@ 1 Introduction
Deep learning methods have become a central driver of progress in medical image analysis, owing to their ability to learn complex hierarchical and semantically rich feature representations from large-scale imaging data. These capabilities enable a broad spectrum of clinical applications, ranging from early disease detection and accurate diagnosis [rana2023machine], malignancy prediction [prakash2020breast, naser2020brain, navarro2021development], and lesion characterization [herent2019detection, yin2023value, dadoun2022deep], to robust image segmentation [ronneberger2015u, isensee2021nnu, cao2022swin, erdur2025deep] and prognostic as well as predictive modeling [zhu2020application, schulz2021multimodal, spohn2023development]. While these capabilities may not only facilitate personalized treatment strategies, they also hold the potential to improve therapeutic effectiveness and patient outcomes [rasuli2025deep]. As advances in imaging hardware continue to accelerate the acquisition of high-resolution tomographic data, the need for automated systems capable of distilling meaningful imaging biomarkers from complex 33D volumes to support clinicians has grown correspondingly [thrall2018artificial]. Effective extraction of such biomarkers is not only essential for alleviating clinician workload but also constitutes a foundational step for enhancing the efficacy of subsequent downstream analytical tasks [rasuli2025deep].
@@
One promising direction to extract semantically rich and versatile imaging features from 22D images lies in the use of vision foundation models, which have demonstrated remarkable generalization capabilities across diverse 22D computer vision tasks [radford2021learning, kirillov2023segment, zou2023segment, oquab2023dinov2, fang2024eva, tschannen2025siglip, venkataramanan2025franca, simeoni2025dinov3]. These models excel at producing both substantially meaningful and robust feature embeddings and can be readily adapted to a wide range of downstream tasks, including segmentation [ayzenberg2024dinov2], classification [veasey2024parameter], or registration [song2024dino]. However, despite their widespread adoption in 22D natural image analysis, efforts to translate foundation model principles to 33D medical imaging have progressed more slowly. A primary constraint remains the absence of large and publicly available multimodal medical imaging datasets required for training at scale [azad2023foundational]. Building on the criteria expressed by paschali2025foundation [paschali2025foundation], a radiology foundation model should couple large-scale architectures with multimodal imaging data, rely on self-supervised training to limit annotation burdens, and exhibit emergent capabilities beyond its explicit training objectives. Yet, despite recent progress and rapid development of large-scale pretrained 33D models, the performance of 33D radiology foundation models still trails that of state-of-the-art 22D vision model counterparts [van2025foundation]. Nevertheless, despite challenges such as the limited availability of large-scale 33D datasets and the computational demands of training 33D models, ongoing research in these areas is likely to lead to substantial breakthroughs and enhanced integration of native 33D models into clinical workflows [azad2023foundational, van2025foundation].
Despite the current performance gap between large-scale 22D and 33D pretrained models, recent studies have demonstrated that 22D vision foundation models trained on natural images, particularly the DINO [caron2021emerging] model family, can be transferred to medical applications with strong results [huix2024natural, liu2025doesdinov3setnew]. Their use, however, necessitates decomposing 33D medical volumes into 22D slices, a process that inevitably disrupts the intrinsic spatial coherence of volumetric data.
@@
A common strategy to combine the resulting slice-wise information extracted by a pretrained encoder is to use a dimension-wise average thereof or a shallow multi-layer perceptron (MLP) [kiechle2024graph, muller2025medical, liu2025doesdinov3setnew]. While this enables end-to-end learning for downstream tasks such as classification, there is no inherent mechanism that explicitly accounts for the relative positions among the slices. More structured approaches attempt to mitigate this limitation: LSTM-based aggregation [donahue2015long, nguyen2020cnn, lilhore2025hybrid] treats sequential slices analogously to temporal data, capturing inter-slice associations, while transformer-based models [muller2025medical, meneghetti2025end] leverage self-attention to more directly model spatial relationships across slices. Still, these methods ultimately remain constrained by the inherent loss of geometry introduced during slice-based decomposition.
@@
With the emergence of graph neural networks (GNNs) [scarselli2008graph], effective modeling of naturally graph-structured data has become possible. GNNs are particularly well-suited for problems involving pairwise interactions or modeling complex relational data by encoding relationships in the graph topology [wu2020comprehensive]. In the context of 33D data, GNNs have demonstrated a strong ability to preserve an underlying spatial structure when the graph topology is derived on the basis of the relative position of 22D representations [Wei_2020_CVPR, Wei_2023_IEEE]. In medical imaging, this makes GNNs well-suited to capture spatial relationships between encoded slices, thereby preserving aspects of the underlying 33D context that are lost in traditional slice-wise representations [kiechle2024graph, di2025structured].
@@
In this work, we introduce TomoGraphView, a novel framework that preserves the spatial structure of volumetric medical images while leveraging the representational power of a 22D foundation model encoder. The name reflects its core components, where Tomo refers to tomography data, Graph denotes the use of GNN-based feature aggregation, and View signifies the integration of multi-view information from several perspectives. More specifically, our framework integrates two key components: omnidirectional volume slicing and spherical graph-based feature aggregation. First, rather than relying solely on canonical axial, coronal, or sagittal planes, we sample cross-sectional views by means of uniformly distributed points on a sphere surrounding the 33D volume, capturing a diverse set of canonical and non-canonical perspectives. These viewpoints are triangulated into a spherical mesh that defines the graph topology, with each node corresponding to features extracted from a 22D vision foundation model applied to a specific slice. Edges between nodes are weighted according to their relative distance, enabling the model to efficiently integrate both local and global context while explicitly preserving spatial structure. By aggregating node features through a GNN, our framework effectively models 3D context from 2D representations, demonstrating the strength of graph-based feature aggregation for volumetric data analysis. The primary contributions of this work are as follows:
@@
(I) We propose a novel out-of-plane slicing technique, termed omnidirectional volume slicing, which extends traditional slicing beyond canonical anatomical axes and yields consistently superior downstream performance compared with standard axial–coronal–sagittal approaches.
@@
(II) We introduce a spherical mesh-based graph topology derived from the omnidirectional sampling strategy, enabling faithful preservation of spatial relationships between multi-view 22D cross-sections within a unified spherical graph-based feature aggregation framework.
@@
(III) Extensive experiments on six oncology 33D medical imaging datasets demonstrate that the proposed TomoGraphView framework, integrating omnidirectional volume slicing with spherical graph-based feature aggregation, consistently outperforms slice-wise aggregation baselines and large-scale pretrained 33D medical models across all datasets and tasks, establishing TomoGraphView as a powerful and generalizable framework for volumetric medical image classification.
@@ 2.1 Slice-based Methods in 33D Medical Imaging
Two-dimensional slice-based approaches remain an appealing alternative to computationally intensive 33D architectures, as they offer substantially lower memory and compute requirements, benefit from the decomposition of 33D volumes into numerous 22D slices, and can leverage powerful pretrained backbones from natural image analysis. These models have repeatedly demonstrated strong performance in medical image classification tasks, underscoring their practical relevance despite the lack of explicit volumetric reasoning [huix2024natural, liu2025doesdinov3setnew].
A substantial body of work highlights the effectiveness of 22D paradigms across diverse clinical settings. navarro2021development [navarro2021development] fine-tuned a DenseNet161 pretrained on ImageNet using axial T1- and T2-weighted MRI slices of soft-tissue sarcomas for tumor grading. Similarly, ahamed2023slice [ahamed2023slice] trained a ResNet18 on axial PET/CT slices of lymphoma to classify whether a slice intersected tumor tissue. wang2024svfr [wang2024svfr] proposed a two-stage slice-to-volume feature representation framework for Alzheimer’s disease, where informative axial MRI slices were selected and aggregated into volume-level features. In parallel, self-supervised foundation models have also been evaluated for slice-level tasks. huang2024comparative [huang2024comparative] and baharoon2024evaluatinggeneralpurposevision [baharoon2024evaluatinggeneralpurposevision] applied DINOv2 to axial brain MRIs and chest CTs for glioma grading and COVID-19 diagnosis, respectively, while liu2025doesdinov3setnew [liu2025doesdinov3setnew] assessed DINOv3 on non-contrast axial CT slices for detecting 18 clinically significant abnormalities.
Building on the limitations of single-plane analysis, several studies have explored multi-plane, or so-called 2.52.5D, strategies that integrate axial, coronal, and sagittal views. saint20222 [saint20222] employed an Xception model trained on 2.52.5D representations to predict HPV status in oropharyngeal cancer. Similarly, meneghetti2025end [meneghetti2025end] explored foundation model-based multiple instance learning for head and neck cancer outcomes, comparing single-plane, multi-plane, and 33D sub-volume representations. A related hybrid approach by jang2022m3t [jang2022m3t] first extracted volumetric features using a 33D CNN, followed by multi-plane slicing of the feature tensor and 22D CNN and transformer processing for Alzheimer’s disease classification.
Despite their empirical success, these methods exclusively rely on canonical anatomical planes aligned with the axial, coronal, and sagittal axes. As a result, structures oriented obliquely or irregularly within the volume may be inadequately represented or omitted entirely. Moreover, beyond the choice of slicing strategy, the mechanism used to aggregate slice-wise features is equally crucial for preserving volumetric context. How these features are combined ultimately determines the extent to which spatial coherence is retained, motivating a deeper examination of aggregation strategies in the subsequent section.
@@ 2.2 Slice-wise Feature Aggregation in 33D Medical Imaging
Slice-wise feature aggregation has emerged as a promising paradigm for 33D medical image analysis, offering a means to leverage powerful 22D foundation models while circumventing the computational burden of end-to-end 33D training. However, because slicing inherently discards volumetric structure, downstream performance depends on how effectively inter-slice spatial dependencies are modeled.
@@
The most straightforward strategy involves collapsing slice embeddings into a single volume-level descriptor using non-learnable operations, such as mean pooling. For example, muller2025medical [muller2025medical] encoded axial slices with DINOv2 and averaged the embeddings for breast MRI, lung CT, and knee MRI classification, while liu2025doesdinov3setnew [liu2025doesdinov3setnew] similarly averaged DINOv3 features from axial chest CT slices for anatomical classification. Although computationally efficient, these approaches ignore inter-slice relationships and therefore provide limited volumetric context.
Learnable fusion mechanisms offer a natural extension. Multi-layer perceptrons (MLPs) have been widely adopted to integrate slice-level features, as in truong2021transferable [truong2021transferable], who fine-tuned self-supervised ImageNet models with an MLP head for several medical classification tasks. baharoon2024evaluatinggeneralpurposevision [baharoon2024evaluatinggeneralpurposevision] scaled this evaluation to 200 radiology experiments across diverse modalities (X-ray, CT, and MRI), and kiechle2024graph [kiechle2024graph] benchmarked MLP aggregation across MedMNIST3D structures. Despite their flexibility, MLPs typically lack explicit awareness of slice order or spatial layout unless augmented with positional conditioning.
Recurrent architectures address this limitation by treating slices as ordered sequences. LSTM-based aggregation has been shown to enhance volumetric reasoning, as demonstrated by zhang2019rsanet [zhang2019rsanet] for multiple sclerosis lesion segmentation and by nguyen2020cnn [nguyen2020cnn] for intracranial hemorrhage detection. Bidirectional variants have also been explored, such as the EfficientNet-B0 + BiLSTM hybrid for breast cancer classification [lilhore2025hybrid]. Slice-by-slice CNN–LSTM hybrids further illustrate the utility of sequential modeling for capturing long-range spatial dependencies [miron2021evaluating].
@@
Transformer-based models provide an alternative that leverages attention to learn global inter-slice relationships. The M3T framework [jang2022m3t] fused multi-plane representations for Alzheimer’s diagnosis, while meneghetti2025end [meneghetti2025end] introduced a foundation model–based multiple instance learning (MIL) transformer for predicting clinical outcomes in head and neck cancer, encoding 22D CT slices with a foundation model and aggregating them using transformer-based attention. More recently, the Medical Slice Transformer (MST) [muller2025medical] combined transformer aggregation with self-supervised 22D encoders, achieving strong results across breast MRI, lung CT, and knee MRI classification downstream tasks.
@@
A complementary direction represents volumetric data as structured graphs. kiechle2024graph [kiechle2024graph] and di2025structured [di2025structured, dipiazza2025structuredspectralgraphrepresentation] modeled tomographic volumes where nodes correspond to slice features and edges encode spatial adjacency. kiechle2024graph [kiechle2024graph] showed that graph representations improve accuracy and runtime over MLP-based aggregation but depend heavily on selecting suitable graph convolution operators and topologies, limiting general applicability. di2025structured [di2025structured, dipiazza2025structuredspectralgraphrepresentation] constructed graphs from axial slice stacks to capture both local and global inter-slice dependencies along the axial axis. However, this design remains restricted to a single orientation, omitting valuable coronal and sagittal context. Moreover, evaluation limited to chest CTs leaves the generalizability of this approach across modalities and anatomical sites uncertain.